OpenStudy (anonymous):

x + 2y = 10 3x + 4y = 8 .Which point is the solution to the system of equations?

OpenStudy (raden):

multiply by 2 the 1st equation then subtract with the 2nd equ

OpenStudy (raden):

2x + 4y = 20 3x + 4y = 8 -------------- (-) -x = 12 x = -12

OpenStudy (raden):

subtitute x=-12 to one of equations above : x + 2y = 10 -12 + 2y = 10 2y = 10 + 12 2y = 22 y = 22/2 = 11 so, as the solution is {-12,11}